A random sample of size n = 598 from a population whose parameter is p = 0.77.
A. What is the mean of the distribution of sample means? Round the answer accurate to 2 decimal places.
B. What is the standard deviation of the distribution of sample means? Round the answer accurate to 2 decimal places.

Answer by Boreal(15235)

Note change:
598*0.77=460.46
variance is np(1-p)=460.46*0.23=105.9058 or 105.91
sd is sqrt(V)=10.29
I rounded only at the end.
(did not see the second "7")